# Ground-Truth Subgraphs for Better Training and Evaluation of Knowledge Graph Augmented LLMs

### v1

@@ Abstract
Retrieval of information from graph-structured knowledge bases represents a promising direction for improving the factuality of LLMs. While various solutions have been proposed, a comparison of methods is difficult due to the lack of challenging QA datasets with ground-truth targets for graph retrieval. We present SynthKGQA [^1] https://github.com/graphcore-research/synth-kgqa, a framework for generating high-quality synthetic Knowledge Graph Question Answering datasets from any Knowledge Graph, providing the full set of ground-truth facts in the KG to reason over each question. We show how, in addition to enabling more informative benchmarking of KG retrievers, the data produced with SynthKGQA also allows us to train better models. We apply SynthKGQA to Wikidata to generate GTSQA [^2] https://huggingface.co/datasets/Graphcore/GTSQA, a new dataset designed to test zero-shot generalization abilities of KG retrievers with respect to unseen graph structures and relation types, and benchmark popular solutions for KG-augmented LLMs on it.
@@ Results
A detailed comparison of the performance of different models on GTSQA is provided in Table˜2. Pure LLMs show limited accuracy with an EM Hits score of at most 33.97, confirming the challenging nature of this new benchmark for models that solely rely on their internal knowledge. Even SOTA KG-RAG models – while performing better than pure LLMs – struggle with the task. Trainable subgraph retrievers outperform KG agents, highlighting the importance of fine-tuning models on the target KG. Among trainable retrievers, SubgraphRAG (with 200 retrieved triples) achieves the best results, surpassing all path-based retrievers, which tend to retrieve smaller subgraphs. The F1 score of retrieved triples is low (<30<30%) across the board, but our results suggest that increasing the recall of ground-truth triples is more beneficial than increasing precision. However, the EM Hits of SubgraphRAG plateaus when increasing the number of retrieved triples further, despite the fact that triple recall continues to improve (Figure˜C7), showing that precision cannot be entirely disregarded. Also note that the recall of ground-truth triples is a much stronger predictor of final model accuracy than the recall of answer nodes (Figure˜C6), proving the value of working with datasets like GTSQA that provide the full ground-truth answer subgraph.
@@ 6 Ground-Truth Subgraphs to Train Better KG Retrievers
While the experiments in the previous paragraphs prove that shortest path triples are not as good a target for retrieval as the ground-truth triples provided by SynthKGQA, it remains to understand whether using them as supervision signal for training subgraph retrievers is equally sub-optimal. We consider three retrievers that use path information for training, namely SR [zhang2022sr], RoG [luo2024rog] and SubgraphRAG [li2024subgraphrag], train them on GTSQA using either the shortest paths or the ground-truth paths between seed and answer nodes as supervision signal, and then compare the performance on the test set. As reported in Table˜3, models trained on the ground-truth subgraphs have EM Hits scores from 55% to 2020% higher than their counterparts trained on shortest paths, due to improved recall (up to +27+27%) and precision (up to +141+141%) of ground-truth triples in the retrieved subgraphs. The improvements are clear for SubgraphRAG (independently of the number of retrieved triples, Figure˜C7), but even more striking for path-based methods. As shown in Figure˜C10, RoG is the model where the quality of the retrieved subgraph benefits more from eliminating the noise in the training data, with precision increasing by more than $9\times$ for 4-hops questions. Similarly, SR experiences a major boost in answer node recall (more than $2\times$ for 4-hops questions). While for 1-hop questions the differences in all statistics are negligible, the gap sharply increases for questions requiring multiple hops (where, as we’ve showed, the shortest paths diverge more appreciably from the ground-truth subgraphs).
@@
Table 3: Improvements in predictive statistics of models trained on ground-truth answer subgraphs (GT), compared to models trained on the shortest paths between seed and answer nodes (SP). For all models, results are averaged over three distinct runs for each of the two training regimes.
@@ C.3 Additional Results
Figure C7: Performance of SubgraphRAG on the test set of GTSQA, as we increase the size of the retrieved subgraph. While recall for both ground-truth triples and answer nodes steadily increases, the final predictive accuracy (EM Hits) eventually plateaus, as too much irrelevant information is added to the LLM prompt. Across all subgraph sizes, training on the ground-truth answer subgraphs (GT) outperforms training on shortest path triples (SP).

### v2

@@ Abstract
Retrieval of information from graph-structured knowledge bases represents a promising direction for improving the factuality of LLMs. While various solutions have been proposed, a comparison of methods is difficult due to the lack of challenging QA datasets with ground-truth targets for graph retrieval. We present SynthKGQA [^1] https://github.com/graphcore-research/synth-kgqa, an LLM-powered framework for generating high-quality Knowledge Graph Question Answering datasets from any Knowledge Graph, providing the full set of ground-truth facts in the KG to reason over questions. We show how, in addition to enabling more informative benchmarking of KG retrievers, the data produced with SynthKGQA also allows us to train better models. We apply SynthKGQA to Wikidata to generate GTSQA [^2] https://huggingface.co/datasets/Graphcore/GTSQA, a new dataset designed to test zero-shot generalization abilities of KG retrievers with respect to unseen graph structures and relation types, and benchmark popular solutions for KG-augmented LLMs on it.
@@ Results
A detailed comparison of the performance of different models on GTSQA is provided in Table˜2. Pure LLMs show limited accuracy with an EM Hits score of at most 33.97, confirming the challenging nature of this new benchmark for models that solely rely on their internal knowledge. Even SOTA KG-RAG models – while performing better than pure LLMs – struggle with the task. Trainable subgraph retrievers outperform KG agents, highlighting the importance of fine-tuning models on the target KG. Among trainable retrievers, SubgraphRAG (with 200 retrieved triples) achieves the best results, surpassing all path-based retrievers, which tend to retrieve smaller subgraphs. The F1 score of retrieved triples is low (<30<30%) across all retrievers, but our results suggest that increasing the recall of ground-truth triples is more beneficial than increasing precision, especially when the final reasoning is performed by more capable LLMs, better at dealing with long context, as we show in the comparison in Figure˜C7. For more compact LLMs, however, retrieval precision should not be entirely disregarded. Also note that the recall of ground-truth triples is a much stronger predictor of final model accuracy than the recall of answer nodes (Figure˜C6), proving the value of working with datasets like GTSQA that provide the full ground-truth answer subgraph.
@@ 6 Ground-Truth Subgraphs to Train Better KG Retrievers
While the experiments in the previous paragraphs prove that shortest path triples are not as good a target for retrieval as the ground-truth triples provided by SynthKGQA, it remains to understand whether using them as supervision signal for training subgraph retrievers is equally sub-optimal. We consider three retrievers that use path information for training, namely SR [zhang2022sr], RoG [luo2024rog] and SubgraphRAG [li2024subgraphrag], train them on GTSQA using either the shortest paths or the ground-truth paths between seed and answer nodes as supervision signal, and then compare the performance on the test set. As reported in Table˜3, models trained on the ground-truth subgraphs have EM Hits scores from 55% to 2020% higher than their counterparts trained on shortest paths, due to improved recall (up to +27+27%) and precision (up to +141+141%) of ground-truth triples in the retrieved subgraphs. The improvements are clear and statistically significant for SubgraphRAG (independently of the number of retrieved triples, Figure˜C7), but even more striking for path-based methods. As shown in Figure˜C10, RoG is the model where the quality of the retrieved subgraph benefits more from eliminating the noise in the training data, with precision increasing by more than $9\times$ for 4-hops questions. Similarly, SR experiences a major boost in answer node recall (more than $2\times$ for 4-hops questions). While for 1-hop questions the differences in all statistics are negligible, the gap sharply increases for questions requiring multiple hops (where, as we’ve showed, the shortest paths diverge more appreciably from the ground-truth subgraphs).
@@
Table 3: Improvements in predictive statistics of models trained on ground-truth answer subgraphs (GT), compared to models trained on the shortest paths between seed and answer nodes (SP). For all models, results are averaged over three distinct runs for each of the two training regimes. We also report (between brackets) the metric’s % variation and the p-value of the null hypothesis that the mean of distribution of SP scores is larger or equal than the one for GT scores (two-sample, one-sided t-test), showing statistical significance.
@@ A.4 Examples
To judge the naturalness and reasonability of questions produced by the SynthKGQA framework, we present three randomly sampled questions for every graph isomorphism type in GTSQA with at least 50 questions in the training set. We observe that the paraphrasing operated by the LLM helps to make the questions sound more natural and human-like.
Example questions and their paraphrased versions Graph isomorphism type (1) Question: On which continent is Palmer Land located?
Paraphrased Question: Which continent is Palmer Land situated on?
Question: What was the military rank of Pierre Gaston-Mayer?
Paraphrased Question: What military rank did Pierre Gaston-Mayer hold?
Question: What was the noble title held by Sir Edward Kerrison, 1st Baronet?
Paraphrased Question: What noble title did Sir Edward Kerrison, 1st Baronet, hold?
Graph isomorphism type (2) Question: Who is the head of government of the capital of the canton of Mérignac-1?
Paraphrased Question: Who is the leader of the government in the capital of the Mérignac-1 canton?
Question: Who is the producer of the series that the episode 1930-talet is part of?
Paraphrased Question: Who produced the series that includes the episode titled "1930-talet"?
Question: Who composed the anthem of Tuvalu?
Paraphrased Question: Who is the composer of Tuvalu’s national anthem?
Graph isomorphism type (3) Question: In which country is the administrative entity in which Arhavi is located, located?
Paraphrased Question: Which country is home to the administrative region where Arhavi is situated?
Question: Where is the mother of the spouse of Ruprecht V of Nassau buried?
Paraphrased Question: Where is Ruprecht V of Nassau’s mother-in-law buried?
Question: Who is the architect of the building occupied by the sports team that Raitis Grafs played for?
Paraphrased Question: Who designed the building where the sports team that Raitis Grafs played for is located?
Graph isomorphism type $(1)$ Question: Which company founded by Henry Herbert Collier has its headquarters in Plumstead?
Paraphrased Question: Which company, established by Henry Herbert Collier, is headquartered in Plumstead?
Question: Which resident of District 2 was killed by Thresh?
Paraphrased Question: Who was the resident of District 2 that Thresh killed?
Question: Which film produced by UK Film Council was based on The Picture of Dorian Gray?
Paraphrased Question: What movie produced by the UK Film Council was inspired by The Picture of Dorian Gray?
Graph isomorphism type (1)$(1)$ Question: Which singer, born in Suphan Buri, holds citizenship of Thailand and plays the guitar?
Paraphrased Question: Which Thai singer from Suphan Buri plays the guitar?
Question: Which film produced by Metro-Goldwyn-Mayer, whose main subject is baseball, was produced by Clarence Brown?
Paraphrased Question: What baseball-themed film made by Metro-Goldwyn-Mayer was directed by Clarence Brown?
Question: Who participated in the 2002 FIFA World Cup and played for both GNK Dinamo Zagreb and Sevilla FC?
Paraphrased Question: Which player took part in the 2002 FIFA World Cup and played for both GNK Dinamo Zagreb and Sevilla FC?
Graph isomorphism type (2)(1) Question: Which team that plays association football is the occupant of the stadium operated by Fenerbahçe Sports Club?
Paraphrased Question: Which football team plays at the stadium managed by Fenerbahçe Sports Club?
Question: What is the city that is both the headquarters location of the developer of Star Wars: Shadows of the Empire (video game) and the narrative location of Sudden Impact?
Paraphrased Question: Which city serves as the headquarters of the developer for the video game Star Wars: Shadows of the Empire and is also the setting for Sudden Impact?
Question: Which street, used for utility cycling, serves as a terminus for a street that is itself a terminus at Frankfurter Tor?
Paraphrased Question: Which street designed for utility cycling acts as a terminus for another street that ends at Frankfurter Tor?
Graph isomorphism type ((1)$(1)$) Question: Which cast member acted in a Spanish-language film that was filmed in Mexico and belongs to the exploitation film genre?
Paraphrased Question: Which cast member starred in a Spanish-language film shot in Mexico that falls under the exploitation genre?
Question: Who is the head of government of the municipality that is contained within Valdizarbe and shares borders with both Uterga and Enériz?
Paraphrased Question: Who is the head of government for the municipality in Valdizarbe that borders both Uterga and Enériz?
Question: Which award was received by the English-language film directed by Terence Davies and starring Freda Dowie?
Paraphrased Question: What award did the English-language film directed by Terence Davies and featuring Freda Dowie win?
Graph isomorphism type (2)$(1)$ Question: Which goalkeeper who participated in the 1958 FIFA World Cup died in a town located in the Eastern Time Zone?
Paraphrased Question: Which goalkeeper from the 1958 FIFA World Cup passed away in a town in the Eastern Time Zone?
Question: Which businessperson who was a cast member of a reality television show also attended the University of Pennsylvania?
Paraphrased Question: Which reality TV star who is also a businessperson went to the University of Pennsylvania?
Question: Which person died of pneumonia in Vienna and has an asteroid belt minor planet named after them?
Paraphrased Question: Who passed away from pneumonia in Vienna and has a minor planet in the asteroid belt named in their honor?
Graph isomorphism type (1)$(1)(1)$ Question: Which musical film directed by Challis Sanderson features both Robb Wilton and Kitty McShane as cast members and has Desmond Dickinson as its director of photography?
Paraphrased Question: What musical film directed by Challis Sanderson stars Robb Wilton and Kitty McShane, with Desmond Dickinson as the director of photography?
Question: Which video game is both an action and platform game, distributed on Nintendo game card for Nintendo DS and published in North America?
Paraphrased Question: What action and platform game is available on a Nintendo game card for the Nintendo DS and was published in North America?
Question: Which minor planet in the asteroid belt was discovered at Osservatorio Astronomico Sormano by both Francesco Manca and Piero Sicoli and is directly preceded by 9110 Choukai?
Paraphrased Question: What is the name of the minor planet in the asteroid belt, discovered by Francesco Manca and Piero Sicoli at the Osservatorio Astronomico Sormano, that comes right before 9110 Choukai?
@@ C.3 Additional Results
Figure C7: Performance of SubgraphRAG (with different LLMs doing the final reasoning) on the test set of GTSQA, as we increase the size of the retrieved subgraph. More capable LLMs are less sensitive on the noise in the retrieved data, with Hits (EM) tracking more closely the increase in recall of ground-truth triples and answer nodes. Across all LLM and subgraph sizes, training on the ground-truth answer subgraphs (GT) outperforms training on shortest path triples (SP).
@@
C.3.1 GPT-5-mini as reasoning model
We repeat the main analysis of Section˜5, but using the newer GPT-5-mini [openai2025gpt5] as the model performing the final reasoning on the KG-retrieved subgraphs, to understand how the choice of the LLM influences Hits (EM). While all models score significantly higher than with GPT-4o-mini (Table˜C3), the relative ranking remains the same as observed before, and no different patterns or behaviors arise when looking at individual graph isomorphism types (Figure˜C11) or at different generalization abilities required to answer (Figure˜C12).
We notice that SubgraphRAG is the model benefiting the most from using a more recent LLM to reason over the retrieved subgraph. This appears to be due to an improved utilization of long context, filtering out the noise in the augmented prompt when presented with a large number of retrieved triples. Indeed, while with GPT-4o-mini we observed a gap between model response quality as measured by Hits (EM) and recall of ground-truth triples when the number of retrieved triples was increased, GPT-5-mini exhibits continued improvements in Hits (EM) up to 500 triples, matching closely the steady increase in GT triple recall (Figure˜C7). This is reflected, across all models, in an even stronger linear correlation between Hits (EM) and the recall of ground-truth triples, with Pearson correlation coefficient increasing from $r=0.95$ ($p=1.3\mathrm{e-}7$) for GPT-4o-mini (Figure˜C6) to $r=0.97$ ($p=1.5\mathrm{e-}8$) for GPT-5-mini. Training on GT triples (rather than shortest paths) proves to be even more beneficial to the final predictive accuracy than we observed for GPT-4o-mini (Figure˜C7).
Table C3: Benchmark of KG-RAG models on GTSQA with GPT-5-mini as final reasoning model.
EM Ground-truth triples Answer nodes
Category Model Hits Recall Recall Precision F1 Hits Recall # triples
KG agent PoG 41.74 40.32 36.07 37.14 34.07 35.76 33.56 3.72
ToG 53.02 51.19 32.78 7.00 10.80 36.13 35.06 8.00
Path-based SR 51.75 49.75 30.22 3.44 5.69 50.25 49.39 72.94
GCR 56.58 54.59 40.71 27.21 29.83 47.10 45.53 6.54
RoG 67.47 65.09 54.69 24.04 27.00 72.91 71.84 72.15
All-at-once SubgraphRAG (200) 78.24 74.28 79.09 1.29 2.53 85.33 84.36 199.61
Figure C11: Hits (EM) of KG-RAG models on different graph isomorphism types, compared to the baseline (GPT-5-mini, no RAG).
Figure C12: Generalization abilities of trainable KG-RAG models. We measure EM performance in terms of difference with the EM of the baseline (GPT-5-mini, no RAG).